OpenStudy (mew55):

line of intersection of the planes 2x+y-z-6=0 and x+4y+2z-1=0

jimthompson5910 (jim_thompson5910):

2x+y-z-6=0 2x+y-6=z z = 2x+y-6 ---------------------- x+4y+2z-1=0 x+4y+2(2x+y-6)-1=0 x+4y+4x+2y-12-1=0 5x+6y-13=0 From here, solve for y (to get the equation in slope-intercept form), but you can stop here or convert to 5x+6y=13 to get the equation in standard form.